OpenStudy (evansanchez):

A 2.0 L flexible container holds 5.0 moles of oxygen (O2) gas. An additional 15.0 moles of nitrogen gas (N2) is added to the container. What is the new volume of the container? 8.0 L 0.13 L 0.50 L 6.0L

OpenStudy (wanell):

Answer 8.0 L. 2.0L / 5.0 moles = x / 20.0 => x = 20 / 5 * 2 = 8